# From Code to Play: Benchmarking Program Search for Games Using Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have shown impressive capabilities in generating program code, opening exciting opportunities for applying program synthesis to games. In this work, we explore the potential of LLMs to directly synthesize usable code for a wide range of gaming applications, focusing on two programming languages, Python and Java. We use an evolutionary hill-climbing algorithm, where the mutations and seeds of the initial programs are controlled by LLMs. For Python, the framework covers various game-related tasks, including five miniature versions of Atari games, ten levels of Baba is You, an environment inspired by Asteroids, and a maze generation task. For Java, the framework contains 12 games from the TAG tabletop games framework. Across 29 tasks, we evaluated 12 language models for Python and 8 for Java. Our findings suggest that the performance of LLMs depends more on the task than on model size. While larger models generate more executable programs, these do not always result in higher-quality solutions but are much more expensive. No model has a clear advantage, although on any specific task, one model may be better. Trying many models on a problem and using the best results across them is more reliable than using just one.
@@ I Introduction
Before the emergence of large language models (LLMs) for code [1], program synthesis in imperative or object-oriented languages like Python or Java was considered highly challenging due to the combinatorial explosion of the search space [2]. Therefore, most solvable tasks were restricted to simple problem domains such as string manipulation or list sorting, typically implemented within a predefined domain-specific language (DSL) [3]. Similarly, program synthesis for games was limited to simple problems with well-defined search spaces, achievable only by incorporating high-level game-specific concepts into the DSL [4, 5, 6, 7].
The use of program synthesis with high-level programming languages in game research has hardly been explored. Most discussions merely outlined its potential applications [8], or focused on the missing aspects of automated game development systems to move from game description languages to programming languages [9].
@@
Our goal is not to propose a new method for program synthesis, but to introduce an easy-to-use and extensible framework to evaluate the current performance of LLMs for the synthesis of game-related program code. To achieve this, we have integrated five different LLM architectures for Python and four for Java, and evaluated 12 and 8 models, respectively. For the synthesis of Python code, the framework consists of five miniature versions of Atari games where the input is represented symbolically [17], ten levels of the game Baba is you, in which various game mechanics are tested [18], a vehicle driving environment based on the game Asteroids, and procedural content generation in the form of mazes. For Java, the framework consists of 12 tabletop games of the TAG framework [16]. In total, we evaluate the LLMs on 29 different tasks. An overview of our proposed framework, as well as games and LLMs used, is shown in Figure 1.
@@ II Related Work
Mariño et al. [7] utilized program search to develop strategies for the game MicroRTS, comparing the resulting programmatic policies with those created by human developers. Their findings demonstrated that the synthesized programs performed comparable to those written by humans. Subsequent research built on this foundation by introducing improved search techniques, including bilevel search [25], by guiding the program search [26] or by searching in semantic spaces [27]. Recently, an approach for combining LLMs with local search algorithms was proposed for MicroRTS [28], where the authors showed that providing initial programs with LLMs found better solutions faster and improved the performance of the final programs.
In [29], Genetic Programming (GP) is used to search for evaluation functions within a predefined DSL for the board game 7 Wonders. This approach resembles our experiments with the TAG framework, where heuristic functions are synthesized; however, we use Java without relying on predefined concepts. GP has also been applied to generate game agents for various scenarios, including a fighting game [30], a platformer game [31], a puzzle game [32], and to create explanations for a maze runner agent [33]. Additionally, Wilson et al. [34] used Cartesian GP to develop programs for Atari games, processing pixel observations through predefined mathematical, statistical, or list functions.
@@ III Framework
The general framework we proposed is based on an evolutionary hill-climbing algorithm where the mutations and the seed of the initial program are performed by an LLM [14, 43]. Thus, our framework belongs to the group of neurosymbolic programming methods [44], as we use an LLM to generate programs that are checked for correctness and functionality by symbolic verifiers, in our case the Python interpreter and Java compiler. The overview of the framework is illustrated in Figure 1, which shows the high-level interaction between the different modules and processes. Synthesized program code by LLMs is always executed within a safe subprocess environment, ensuring that the main process can terminate it after a certain time limit to prevent infinite execution of the program.
A detailed description of the complete algorithm is provided in Algorithm 1. Our framework consists of two iterative processes that control the length of the search, defined by the input parameter iterations, as well as the number of attempts to generate, repair or improve the program in each iteration defined by the input parameter maxAttempts. Each iteration starts by generating a task prompt to obtain an initial Python or Java function, and inject and run the code in a subprocess. Afterwards, the inner iterative process starts, where the task prompt is updated and the program is repaired or improved. If the function is executed successfully, we update the prompt with the achieved evaluation metric and all relevant environment-specific details, such as the action trace of the executed function. If an error occurs, e.g. a syntax problem, a runtime error due to improper array indexing or similar problems, the error description is included in the prompt for refinement of the program. These steps are repeated iteratively until the evaluation criteria defined by the fitness function are satisfied or the specified number of maxAttempts is reached.
The outer loop is able to stop the current program search and restart from the initial prompt or the last successful program found. In order to achieve this we introduce the variable blankRestart, which is an input parameter to Algorithm 1. This is necessary when the LLM tries to fix compilation errors in the case of Java or goes astray, such as when generating DQN code for the Atari environments, which we experienced in the preliminary experiments. This also depends on the problem domain, therefore blankRestart is adapted to the corresponding domain.
@@
Algorithm 1 The algorithm for our proposed framework, which consists of two iterative processes that control the length of the search and the number of attempts to generate, repair or improve the program in each iteration.
Input: task, iterations, maxAttempts, blankRestart
      Output: program
1:procedure ProgramSearch
2:     Init Vars: lastProgram, lastResult, lastFitness $\leftarrow$ 0
3:     for $i\leftarrow 0$ to iterations do
4:         if blankRestart or lastFitness ==absent$=$ =then
5:              prompt $\leftarrow$ GetTaskPrompt(task)
6:         else
7:              prompt $\leftarrow$ UpdateTaskPrompt(task, lastResult, lastFitness, lastProgram)
8:         end if
9:         program $\leftarrow$ QueryLlm(prompt)
10:         result $\leftarrow$ InjectAndRunCode(task, program)
11:         fitness $\leftarrow$ EvaluateFitness(task, result)
12:         $j\leftarrow 1$
13:         while not CheckCriterion(fitness) or
14:     $j<$ maxAttempts do
15:              prompt $\leftarrow$ UpdateTaskPrompt(task, result, fitness, program)
16:              program $\leftarrow$ QueryLlm(prompt)
17:              result $\leftarrow$ InjectAndRunCode(task, program)
18:              fitness $\leftarrow$ EvaluateFitness(task, result)
19:              $j\leftarrow j+1$
20:              if fitness $>$ lastFitness then
21:                  lastProgram $\leftarrow$ program
22:                  lastResult $\leftarrow$ result
23:                  lastFitness $\leftarrow$ fitness
24:              end if
25:         end while
26:     end for
27:return program
28:end procedure
@@ IV Large Language Models
LLM Python Java
Llama 3.1 8B 2024-04-18 -
Llama 3.1 70B 2024-04-18 -
Llama 3.1 405B 2024-04-18 -
Claude 3.5 Haiku 2024-10-22 2024-03-07
Claude 3.5 Sonnet 2024-10-22 2024-06-20
GPT 4o mini 2024-07-18 2024-07-18
GPT 4o 2024-08-06 2024-08-06
o1-mini 2024-09-12 -
Mistral Small 2024-09 2024-09
Mistral Large 2024-07 2024-07
Gemini Pro 2024-09 2024-09
Gemini Flash 2024-09 2024-09
@@
For our benchmark, we integrated five LLM providers for Python and four for Java in our framework, namely the small and the large version for each model type. From OpenAI, we utilize models from the GPT-4o family [^3] https://platform.openai.com/docs/models, based on GPT-4 [45]. We also incorporate the latest models from Mistral [^4] https://docs.mistral.ai/getting-started/models/models_overview/, Claude 3.5555https://docs.anthropic.com/en/docs/about-claude/models from Anthropic, based on Claude 3 [46], and the latest Gemini models [^6] https://ai.google.dev/gemini-api/docs/models/gemini [47], provided by Google, for both programming languages.
Models from the Llama 3.1 family [48] are included only for Python tasks, as they are not supported by the LangChain4j library, [^7] https://docs.langchain4j.dev which we integrated into the TAG framework. For the new ChatGPT models in the o1 generation, we use o1-mini, which offers performance comparable to o1-preview for coding tasks. [^8] https://openai.com/index/openai-o1-mini-advancing-cost-efficient-reasoning/ However, this model is also limited to Python.
Details on the model versions and their release dates are summarized in Table I.
@@ V Game Applications
In the following section, we describe the experiments that were conducted for each of our target domains.
V-A Programmatic Policies: Minatar
Minatar [17] is a collection of five games that are miniature versions of Atari games. In Minatar, the games are represented as a symbolic state space on a 10$\times$10×$\times n$ grid, where $n$ represents the number of channels, and each channel represents an object such as walls, enemies or the agent. Minatar is an ideal test bed for experiments, as the games are more efficient to learn without changing the game mechanics of the original game. Previously, Minatar was used in [24] to explain the behaviour of agents through program synthesis, but it was only possible to explain short sub-trajectories since enumerative search-based methods were used to search through a predefined domain-specific language that resembles Lisp. In our experiments, we use all available Minatar environments, which are shown in Figure 2. The game descriptions are outlined below:
  • Seaquest: In this game, the agent controls a submarine and is able to shoot bullets. The objective is to save as many divers as possible, while also shooting enemy submarines or sharks. Each time an enemy is struck, the reward is increased by one. When the submarine saves the divers, the agent also receives a reward.
@@ V-A Programmatic Policies: Minatar
  • Asterix: The objective of the game is to collect gold while avoiding enemies. The player gets one reward for each collected gold and the game is over when the player is hit by an enemy.
  • Space Invaders: The agent controls a cannon and shoots aliens while dodging bullets launched from the alien spaceship. Additionally, the player must prevent the aliens from reaching the bottom of the screen. For each destroyed alien, one reward is received.
  • Breakout: The goal is to destroy all the bricks with the ball by controlling the paddle to bounce the ball off before it goes out off the screen. With each destroyed brick the agent receives a reward of one.
@@
The LLMs were prompted to generate a Python function which can be used as a policy to play the game. The prompt contains information about the game rules, the objective of the agent and also the possible actions of the environment and available game objects. The description of the game, was taken from Young and Tian [17]. The prompts for the games are available in the code repository [^9] https://github.com/ManuelEberhardinger/Benchmarking-Language-Model-Based-Program-Search-for-Games. The LLM receives only the initial state, which is preprocessed from the state input of the environment, a one-hot encoded 3D array, into a 2D array with text descriptions for each grid cell representing the cell’s object. Figure 3 shows an example of the converted state for Breakout. All other games convert the state in a similar way so that the LLM can process the state input semantically.
Each of the games tests the LLM for different game concepts. Space Invaders, Breakout and Freeway restrict the agent’s movement by only allowing horizontal or vertical movement. Space Invader and Seaquest allow the player to fight the enemy, while in Asterix and Freeway the player can only avoid the enemies. In Asterix, the player must also collect items in order to receive a reward. Seaquest is the most difficult game, as the player has to collect six divers and then reach the surface so that the divers can leave the submarine, but at the same time the player has to shoot down enemies. Breakout, on the other hand, is one of the easier games as there are no opponents and the player only has to anticipate where the ball will land in order to bounce it off with the paddle.
@@
TABLE II: Average reward of all executable programs for the Minatar experiments with the successful programs for each game. In each case, 10 iterations of the search were performed, using 3 queries in each iteration to create or improve the policy. For each program, 50 evaluation episodes were performed.
@@
Llama 3.1 8B 0.0 (5) 0.0 (7) 0.27 (7) 0.49 (4) 1.5 (6)
Llama 3.1 70B 0.07 (10) 1.88 (10) 1.75 (10) 1.72 (9) 3.32 (10)
Llama 3.1 405B 0.12 (10) 1.15 (10) 1.92 (10) 2.51 (9) 5.79 (10)
Claude 3.5 Haiku 0.18 (10) 1.96 (10) 2.99 (10) 2.81 (10) 6.6 (10)
Claude 3.5 Sonnet 0.72 (10) 0.56 (10) 4.16 (10) 2.92 (10) 11.59 (10)
GPT 4o mini 0.06 (10) 1.18 (9) 0.82 (10) 1.49 (10) 4.62 (10)
GPT 4o 0.17 (10) 4.23 (10) 2.73 (10) 3.88 (10) 6.26 (10)
o1-Mini 0.62 (10) 3.94 (10) 2.65 (10) 3.12 (10) 9.48 (10)
Mistral Small 0.0 (0) 1.16 (5) 0.22 (4) 1.02 (3) 4.31 (8)
Mistral Large 0.0 (10) 5.25 (10) 3.72 (10) 2.08 (8) 5.21 (10)
Gemini Flash 0.0 (10) 1.19 (7) 1.21 (9) 1.62 (7) 3.77 (10)
Gemini Pro 0.08 (10) 1.72 (10) 2.87 (10) 1.81 (6) 4.44 (10)
For all Minatar experiments, we use 10 iterations with max three attempts in each iteration, which results in max 30 prompts for each model. For each program, 50 evaluation episodes were performed. Table II shows the average reward of all executable programs for each LLM with the number of successful iterations in the brackets – an executable program that returns a positive reward. In general, the larger models perform better than their smaller counterparts in terms of average reward, with Claude Sonnet being the best performing model. Only at Freeway the Claude Haiku and Llama 70B models beat their larger counterparts in the family. o1-mini also outperforms GPT 4o in Space Invader and Seaquest. This is, however, not represented by the maximum reward shown in Table III. Only in Breakout, Claude Sonnet is the top-performing model regarding the average and maximum reward, but is beaten in three games in terms of the maximum reward by Claude Haiku. This pattern is also visible in the other LLM families where the small model outperforms their larger version.
It can also be seen from both tables that it is more difficult to find good programs for more complicated games such as Seaquest. Only o1-mini, which is praised by OpenAI for its sophisticated reasoning capabilities, performs well in this game, but fails to beat the other LLMs on the simpler games. Looking at the best programs for each LLM, o1-mini manages to correctly locate enemies and shoot them if they are in a line, while the other programs only check if enemies are nearby without checking if they are in a line. o1-mini is also the only model that uses the Manhattan distance to move to nearby enemies, while all other programs only try to shoot enemies or rescue divers.
For Freeway, all of the best performing programs show similar behavior for each LLM, while the poor performing models struggle to correctly implement a one step lookahead (OSLA) of the cars. The best performing program of Mistral Large is the only program that uses the modulo operation to correctly predict when the car will be teleported to the other side when it drives out of the screen, which was mentioned in the prompt.
For Breakout, the best programs of Mistral Small and Llama 8B only receive points by chance, as they return a random action or do not take the position of the ball into account. All other programs are able to locate the ball and also the direction in which the ball is moving and move the paddle in that direction. Claude Sonnet is the only model that correctly uses OSLA to predict the next column of the ball. GPT 4o also uses OSLA, but confuses the row with the column.
In Asterix, the best performing model Claude Haiku prioritizes the gold, but also checks whether it is moving towards an enemy as it approaches the gold. The mediocre performing models often prioritize the gold without checking if there are enemies nearby. o1-mini, instead, uses action values that are updated depending on nearby gold coins and enemies, but is still not able to beat Claude Haiku.
For Space Invader, the good performing programs with a reward over 20, correctly locate enemy bullets, aliens and the cannon. The programs below 20, do confuse sometimes the column with the row or only take into account the first alien found in the state. The smallest LLM, Llama 8B, produces a program that fires when there are no friendly bullets, i.e. bullets shot by the program itself, in the state, and was still able to get a reward of 5.9 even though the program makes no sense at all.
Overall, it can be said that in the Minatar games larger models in most cases show a more sophisticated behavior in the programs, but as can be seen with Claude Haiku, this is not always the case. Currently, only a very simple prompting strategy is used, which already gives good results in most games. Using more complicated prompting strategies, such as Chain of Thought [49] or adding a crossover operator could lead to improvements in the programs found.
TABLE III: Max Reward of the best program for the Minatar experiments with the same experiment setup as Table II.
Model Seaquest Freeway Breakout Asterix SpaceInvader
Llama 3.1 8B 0.0 0.0 1.1 1.2 5.9
Llama 3.1 70B 0.6 8.7 9.3 4.4 6.3
Llama 3.1 405B 0.6 4.8 9.0 6.4 20.2
Claude 3.5 Haiku 0.8 7.5 12.2 11.0 24.5
Claude 3.5 Sonnet 2.3 5.3 20.2 6.9 22.8
GPT 4o mini 0.7 5.6 3.0 3.8 17.8
GPT 4o 1.1 9.5 7.3 8.7 22.4
o1-mini 11.3 8.8 4.8 10.4 22.6
Mistral Small 0.0 5.8 0.7 3.9 21.9
Mistral Large 0.0 9.9 7.7 5.9 11.9
Gemini Flash 0.0 7.4 5.6 3.0 13.3
Gemini Pro 0.8 6.4 6.2 4.8 15.4
V-B Vehicle Driving
@@
TABLE IV: The minimum distance of the best program for the Asteroids ship driving experiments for different rotation speeds $\omega$ in degrees per second. In each case, 10 iterations of the search were performed, using 3 queries in each iteration to create or improve the policy. For each program, the same set of 5 evaluation tasks were used, each with different target positions and initial states of the ship.

### v2

@@ Abstract
Large language models (LLMs) have shown impressive capabilities in generating program code, opening exciting opportunities for applying program synthesis to games. In this work, we explore the potential of LLMs to write usable code for a wide range of gaming applications, focusing on two programming languages, Python and Java. We use a hill-climbing algorithm, where the mutations and seeds of the initial programs are controlled by LLMs. For Python, the framework covers various game-related tasks, including five miniature versions of Atari games, ten levels of Baba is You, an environment inspired by Asteroids, and a maze generation task. For Java, the framework contains 12 games from the TAG tabletop games framework. Across 29 tasks, we evaluated 11 language models for generating Python and Java code. Our findings suggest that the performance of LLMs depends more on the task than on model size. In addition, the experiments show that running the program search algorithm multiple times with fewer iterations is better than a single run with more iterations.
@@ I Introduction
Before the emergence of large language models (LLMs) for code [1], program synthesis in imperative or object-oriented languages like Python or Java was considered a highly challenging task due to the combinatorial explosion of the search space [2]. Therefore, most solvable tasks were restricted to simple problem domains such as string manipulation or list sorting, typically implemented within a predefined domain-specific language (DSL) [3]. Similarly, program synthesis for games was limited to simple problems with well-defined search spaces, achievable only by incorporating high-level game-specific concepts into the DSL [4, 5, 6, 7].
The goal of program synthesis is the automatic generation of programs that satisfy certain requirements, such as input-output examples (e.g. five pairs of unsorted lists and their sorted counterparts), natural language descriptions of the program to be generated or formal specifications such as logical constraints [2]. Automatic construction of programs from high-level programming languages in game research has hardly been explored. Most discussions merely outlined its potential applications [8], or focused on the missing aspects of automated game development systems to move from game description languages to programming languages [9].
@@
Our goal is not to propose a new method for program synthesis, but to introduce an easy-to-use and extensible framework to evaluate the current performance of LLMs for the synthesis of game-related program code. To achieve this, we have integrated five different LLM providers with a total of 11 different models. For the synthesis of Python code, the framework consists of five miniature versions of Atari games where the input is represented symbolically [17], ten levels of the game Baba is You, in which various game mechanics are tested [18], a vehicle driving environment based on the game Asteroids, and procedural content generation in the form of mazes. For Java, the framework consists of 12 tabletop games of the TAG framework [16]. In total, we evaluate the LLMs on 29 different tasks. An overview of our proposed framework, as well as games and LLMs used, is shown in Figure 1.
@@
The structure of the paper is as follows: We review related work on program synthesis for games in Section II. Afterwards, the framework is introduced in Section III, followed by a short summary of the LLMs used in Section IV. Section V first describes the experimental setup (Section V-A) and then the experiments for the game applications, starting with miniature versions of the Atari games, followed by the Asteriods-inspired vehicle driving experiments. Then we discuss the puzzle game Baba is You is and the maze generation experiments. We conclude the first part of the experiments with an overall evaluation of the generated Python code. The second part of the experiments, starting in Section V-G, focuses on the TAG framework which uses Java as the programming language. In addition to the standardised experiments, longer-running experiments that evaluate the impact on the number of iterations for the five games in the Atari domain are reported in Section VI. Finally, we discuss practical insights gained from the experiments in Section VII, and conclude the paper in Section VIII.
@@ II Related Work
Mariño et al. [7] utilized program search to develop strategies for the game MicroRTS, comparing the resulting programmatic policies with those created by human developers. Their findings demonstrated that the synthesized programs performed comparably to those written by humans. Subsequent research built on this foundation by introducing improved search techniques, including bilevel search [25], by guiding the program search [26] or by searching in semantic spaces [27]. Recently, an approach for combining LLMs with local search algorithms was proposed for MicroRTS [28], where the authors showed that providing initial programs with LLMs found better solutions faster and improved the performance of the final programs.
In [29], Genetic Programming (GP) is used to search for evaluation functions within a predefined DSL for the board game 7 Wonders. Similarly, Sturtevant and White generate evaluation functions automatically by using reinforcement learning to learn more complex features by combining atomic features for the game Hearts [30]. This approach resembles our experiments with the TAG framework, where heuristic functions are synthesized; however, we use Java without relying on predefined concepts. GP has also been applied to generate game agents for various scenarios, including a fighting game [31], a platformer game [32], a puzzle game [33], and to create explanations for a maze runner agent [34]. Additionally, Wilson et al. [35] used Cartesian GP to develop programs for Atari games, processing pixel observations through predefined mathematical, statistical, or list functions.
@@ III Framework
The general framework we propose is based on a hill-climbing algorithm where the mutations and the seed of the initial program are performed by an LLM [14, 44]. Thus, our framework belongs to the group of neurosymbolic programming methods [45], as we use an LLM to generate programs that are checked for correctness and functionality by symbolic verifiers, in our case the Python interpreter and Java compiler. The overview of the framework is illustrated in Figure 1, which shows the high-level interaction between the different modules and processes. Synthesized program code by LLMs is always executed within a safe subprocess environment, ensuring that the main process can terminate it after a certain time limit to prevent infinite execution of the program.
A detailed description of the complete algorithm is provided in Algorithm 2. Our framework consists of two iterative processes that control the length of the search, defined by the input parameter iterations, as well as the number of attempts to repair the program in each iteration defined by the input parameter maxRepairs. Each iteration starts by generating or updating a task prompt to obtain an initial Python or Java function from Algorithm 1. The procedure of Algorithm 1 queries the LLM and tries to fix compilation or syntax errors for the given number of maxRepairs. The program code is then returned to Algorithm 2 and is injected and executed in a subprocess. If the function is executed successfully and returns an improved reward, we update the prompt with the achieved evaluation metric and all relevant environment-specific details, such as the action trace of the executed function. If a runtime error occurs, the error description is included in the prompt for refinement of the program, so that it is possible for the LLM to fix these errors in the next iteration. These steps are repeated iteratively until the evaluation criteria defined by the fitness function are satisfied or the specified number of iterations is reached.
@@
Algorithm 1 The algorithm for querying an LLM with the number of attempts to generate/update or repair the program.
1:procedure QueryLlmWithRepair(prompt, maxRepairs)
2:     program $\leftarrow$ QueryLlm(prompt)
3:     $j\leftarrow 1$
4:     while $j<$ maxRepairs and not CanExec(program) do
5:         prompt $\leftarrow$ GetRepairPrompt(program)
6:         program $\leftarrow$ QueryLlm(prompt)
7:         $j\leftarrow j+1$
8:     end while
9:return program
10:end procedure
Algorithm 2 The algorithm for our proposed framework, which uses the procedure from Algorithm 1 in each iteration to improve or generate a program.
1:procedure ProgramSearch(task, iterations, maxRepairs)
2:     bestFitness $\leftarrow$ 0
3:     $i\leftarrow 1$
4:     repeat
5:         if bestProgram then
6:              prompt $\leftarrow$ UpdateTaskPrompt(task, bestResult, bestFitness, bestProgram)
7:         else
8:              prompt $\leftarrow$ GetTaskPrompt(task)
9:         end if
10:         program $\leftarrow$ QueryLlmWithRepair(prompt, maxRepairs)
11:         result $\leftarrow$ InjectAndRunCode(task, program)
12:         fitness $\leftarrow$ EvaluateFitness(task, result)
13:         if fitness $>$ bestFitness then
14:              bestProgram $\leftarrow$ program
15:              bestResult $\leftarrow$ result
16:              bestFitness $\leftarrow$ fitness
17:         end if
18:         $i\leftarrow i+1$
19:     until not CheckCriterion(fitness) and $i<$ iterations
20:return program
21:end procedure
@@ IV Large Language Models
LLM Release Date
Claude 3.5 Sonnet 2024-10-22
Claude 3.5 Haiku 2024-10-22
Gemini 2.0 Flash 2025-02
Gemini 2.0 Flash-Lite 2025-02
Mistral Large 2024-11
Mistral Small 2025-03
o1-mini 2024-09-12
GPT 4o 2024-08-06
GPT 4o mini 2024-07-18
Llama 3.3 70B 2024-12-06
Llama 3.1 70B 2024-07-23
For our benchmark, we integrated five LLM providers for generating Python and Java code in our framework, using one smaller and one larger version for each model type. From OpenAI, we utilize models from the GPT-4o family [^3] https://platform.openai.com/docs/models, based on GPT-4 [46]. For the new ChatGPT models in the o1 generation, we use o1-mini, which offers performance comparable to o1-preview for coding tasks. [^4] https://openai.com/index/openai-o1-mini-advancing-cost-efficient-reasoning/ We also incorporate the latest models from Mistral [^5] https://docs.mistral.ai/getting-started/models/models_overview/, Claude 3.5666https://docs.anthropic.com/en/docs/about-claude/models from Anthropic, based on Claude 3 [47], and the 2.0 Gemini Flash models [^7] https://ai.google.dev/gemini-api/docs/models/gemini [48], provided by Google, for both programming languages. Two open source models from the Llama family are used, Llama 3.1 70B [49] and Llama 3.3 70B, which performs similarly to Llama 3.1 405B for code-related tasks. [^8] https://www.llama.com/docs/model-cards-and-prompt-formats/llama3_3/ Details on the model versions and their release dates are in Table I.
@@ V Game Applications
In the following section, we first describe the experimental setup, followed by the experiments that were conducted for each of our target domains.
V-A Experimental Setup
In general, the experiments reported run 10 independent trials of Algorithm 2, and the final recommendation is the best performing agent from these 10 trials. In each trial, 10 iterations of search were performed, with 3 queries used to create/improve or repair the policy.
V-B Programmatic Policies: Minatar
Minatar [17] is a collection of five games that are miniature versions of Atari games. In Minatar, the games are represented as a symbolic state space on a 10$\times$10×$\times n$ grid, where $n$ represents the number of channels, and each channel represents an object such as walls, enemies or the agent. Minatar is an ideal test bed for experiments, as the games are more efficient to learn without changing the game mechanics of the original game. Previously, Minatar was used in [24] to explain the behaviour of agents through program synthesis, but it was only possible to explain short sub-trajectories since enumerative search-based methods were used to search through a predefined domain-specific language that resembles Lisp. In our experiments, we use all available Minatar environments, which are shown in Figure 2. The games are:
  • Seaquest: The agent controls a submarine and is able to shoot bullets. The objective is to save as many divers as possible, while also shooting enemy submarines or sharks. Each time an enemy is struck, the reward is increased by one. When the submarine saves the divers, the agent also receives a reward.
@@ V-B Programmatic Policies: Minatar
  • Asterix: The objective of the game is to collect gold while avoiding enemies. The player gets one point for each collected gold and the game is over when the player is hit by an enemy.
  • Space Invaders: The agent controls a cannon and shoots aliens while dodging bullets launched from the alien spaceship. Additionally, the player must prevent the aliens from reaching the bottom of the screen. For each destroyed alien, one point is received.
  • Breakout: The goal is to destroy all the bricks with the ball by controlling the paddle to bounce the ball off before it goes out off the screen. With each destroyed brick the agent receives one point.
@@
The LLMs were prompted to generate a Python function which can be used as a policy to play the game. The prompt contains information about the game rules, the objective of the agent and also the possible actions of the environment and available game objects. The description of the game was taken from Young and Tian [17]. The prompts for the games are available in the code repository [^9] https://github.com/ManuelEberhardinger/Benchmarking-Language-Model-Based-Program-Search-for-Games. The LLM receives only the initial state, which is preprocessed from the state input of the environment, a one-hot encoded 3D array, into a 2D array with text descriptions for each grid cell representing the cell’s object. Figure 3 shows an example of the converted state for Breakout. All other games convert the state in a similar way so that the LLM can process the state input semantically.
Each of the games tests the LLM for different game concepts. Space Invaders, Breakout and Freeway restrict the agent’s movement by only allowing horizontal or vertical movement. Space Invader and Seaquest allow the player to fight the enemy, while in Asterix and Freeway the player can only avoid the enemies. In Asterix, the player must also collect items in order to receive a reward. Seaquest is the most difficult game, as the player has to collect six divers and then reach the surface so that the divers can leave the submarine, but at the same time the player has to shoot down enemies. Breakout is one of the easier games compared to the others, as there are no opponents and the player only has to anticipate where the ball will land in order to bounce it off with the paddle.
TABLE II: Max Reward of the best program for the Minatar experiments with 50 evaluation episodes.
Model Seaquest Freeway Breakout Asterix SpaceInvader
Claude Sonnet 2.43 7.22 18.43 11.37 29.07
Claude Haiku 3.68 7.33 7.13 6.78 21.91
Gemini Flash 2.98 8.07 12.3 6.09 26.34
Gemini Lite 0.58 6.49 15.56 5.06 16.57
Mistral Large 1.3 8.48 12.03 6.48 20.89
Mistral Small 0.26 8.2 5.59 5.88 22.41
o1 mini 5.73 9.52 21.81 10.39 22.45
GPT 4o 1.15 10.05 15.26 10.91 23.83
GPT 4o mini 0.47 8.44 8.56 6.55 27.45
Llama 3.3 70B 4.01 7.74 8.85 5.8 17.1
Llama 3.1 70B 4.08 8.02 4.39 5.54 13.04
@@
TABLE III: Average reward of the 10 programs found after each trial. We also report the standard deviation $\sigma$ of the 10 programs after each trial. The min. reward for all games is 0, which indicates that there is a high variance in between the trials.
@@
Claude Sonnet 1.56$\pm$0.38 3.32$\pm$2.78 8.25$\pm$4.1 6.52$\pm$2.46 22.38$\pm$5.04
Claude Haiku 1.07$\pm$0.96 3.8$\pm$3.03 3.77$\pm$1.85 5.56$\pm$0.85 14.5$\pm$5.53
Gemini Flash 1.04$\pm$0.75 3.31$\pm$2.83 5.82$\pm$3.34 4.36$\pm$1.22 9.98$\pm$9.27
Gemini Lite 0.14$\pm$0.19 0.65$\pm$1.95 6.8$\pm$5.25 3.13$\pm$1.69 5.02$\pm$4.12
Mistral Large 0.29$\pm$0.45 6.6$\pm$2.3 5.69$\pm$2.67 4.75$\pm$1.12 12.62$\pm$4.85
Mistral Small 0.07$\pm$0.09 3.44$\pm$3.15 2.96$\pm$1.97 3.51$\pm$1.67 9.46$\pm$6.76
o1 mini 1.49$\pm$1.56 7.9$\pm$1.81 8.58$\pm$4.99 7.53$\pm$1.46 19.48$\pm$3.56
GPT 4o 0.36$\pm$0.36 7.75$\pm$1.56 6.6$\pm$3.97 7.36$\pm$2.85 15.29$\pm$4.7
GPT 4o mini 0.16$\pm$0.18 3.94$\pm$3.97 3.53$\pm$2.47 3.88$\pm$1.98 6.95$\pm$7.85
Llama 3.3 70B 0.79$\pm$1.14 3.22$\pm$2.26 3.42$\pm$2.86 3.87$\pm$1.39 6.16$\pm$4.07
Llama 3.1 70B 0.7$\pm$1.17 5.23$\pm$2.51 2.06$\pm$1.33 3.03$\pm$1.58 5.76$\pm$3.9
For each program, 50 evaluation episodes were performed. All trials resulted in successful programs, i.e., programs that were executable and returned a positive reward. Table II presents the max. reward of the best program found and Table III shows the average reward with the standard deviation of the 10 programs found in each trial.
In general, larger models outperform their smaller counterparts in average reward, with o1 mini and Claude Sonnet achieving the highest overall performance. This trend is also reflected in the maximum reward: Claude Sonnet generated the best programs for Space Invaders and Asterix, while o1 mini produced the best results for Seaquest and Breakout. GPT 4o found the best performing program for Freeway, although o1 mini slightly outperforms it on average.
In the case of Space Invaders, GPT 4o mini achieved the second best program, though this appears to be a special case, since its average reward is approximately three times lower than that of Claude Sonnet, suggesting that the high performance may be due to chance. The standard deviation across the top 10 programs is often quite high, indicating that multiple trials of program search may be more effective than a single long run. To explore this hypothesis, we conduct a long running experiment in the Minatar domain, presented in section VI, where we also show reward curves to determine when the best programs is found. Interestingly, in some cases smaller models outperform their larger counterparts, e.g., Claude Haiku in Seaquest and Freeway and Gemini Flash Lite in Breakout. It can be seen from both tables that it is more difficult to find good programs for more complicated games such as Seaquest.
Looking at the best programs for each LLM in Seaquest, o1 mini is the only model that defines multiple utility functions to check if it can shoot enemies, or implements one-step look-ahead (OSLA) to check if nearby cells are safe to move to.
In Freeway, all the best performing programs show similar behaviour for each LLM and implement some form of OSLA to check whether it is safe for the chicken to cross the road. GPT 4o’s best performing program even uses three look-ahead steps, i.e. it only moves forward if the three cells above the chicken are empty.
@@
TABLE IV: The minimum distance of the best program for the Asteroids ship driving experiments for different rotation speeds $\omega$ in degrees per second. $D_{avg}$ shows the average distance of the best programs for the different rotation speeds. For each program, the same set of 5 evaluation tasks were used, each with different target positions and initial states of the ship. The heatmap uses an inverted colour gradient, i.e., values close to zero are darker.